OpenStudy (anonymous):

rationalize denominator of 3/2+ the square root of 3

OpenStudy (campbell_st):

if this is the question you will need to multiply by 1 \[\frac{3}{2+\sqrt{3}}\] by 1 is written as \[\frac{2 - \sqrt{3}}{2 - \sqrt{3}}\] so you will get a rational number in the denominator \[\frac{3}{2 + \sqrt{3}} \times \frac{ 2 - \sqrt{3}}{2 - \sqrt{3}}\]

OpenStudy (anonymous):

thats why i got to when i did the problem then i could not figure out the solution. do u know how

OpenStudy (anonymous):

what*

Directrix (directrix):

@lexigolds Calculate the product of the numerators and the product of the denominators as @campbell_st has indicated. To help you get started, the product of the numerators would be 3( 2 - sqrt(3) ) = 6 - 3*sqrt(3) (Use the Distributive Property). Use multiplication of binomials technique with the denominators.